OpenStudy (naga447):

I don't know what to do can someone please explain how I solve this.... Solve for x: 3|x - 3| + 2 = 14

OpenStudy (zzr0ck3r):

can you get it so that we have |stuff| = other stuff

OpenStudy (zzr0ck3r):

so get |x-3| by itself

OpenStudy (naga447):

Yeah I understand absolute values I just don't have any idea what i'm trying to do in this equation...

OpenStudy (zzr0ck3r):

again, can you get |x-3| by itself @naga447 ?

OpenStudy (anonymous):

@Hgskimh ask what you want..

OpenStudy (zzr0ck3r):

@Hgskimh dont ask what you want. This is not your question and its rude.

OpenStudy (naga447):

3|x - 3| + 2 = 14 3|x - 3| = 12 |x - 3| = 12 / 3 |x - 3| = 4 so what is the next step?

OpenStudy (zzr0ck3r):

ok good

OpenStudy (zzr0ck3r):

now we have this thing that says \(|a|=b \iff a=b \text{ or } a=-b\) so your solutions are given by the two equations \(x-3=4\) and \(x-3=-4\) solve them and you are done.

OpenStudy (naga447):

Okay thanks!

OpenStudy (zzr0ck3r):

no pee z